# Towards More Accurate US Presidential Election via Multi-step Reasoning with Large Language Models

### v2

@@ Abstract
Important Notice Ongoing Work: This research is ongoing as of November 27, 2024. Research Integrity: This research is conducted independently w/o any funding. Content Warning: This paper may contain some offensive content generated by LLMs. Disclaimer: This study explores LLMs’ capacity in election forecasting. Predictions do not reflect the authors’ views and should not be interpreted as definitive forecasts.
@@ 2.1 LLMs in Political Science: A New and Emerging Field
2.2 Political Election Research: Classical and Agent-Based Approaches
Traditional political science literature has long relied on survey data and statistical models to analyze voter behavior and predict election outcomes. Classical models like the Downsian spatial model and the Median Voter Theorem explain election results by assuming voters’ positions in policy space Downs (1957); Black (1948). Time-series regression models also play a key role in long-term election forecasting, analyzing the relationships between economic indicators, political events, and voter sentiment Lewis-Beck and Rice (1992); Erikson and Wlezien (2016).
In recent years, Agent-Based Models (ABMs) have gained traction in election research. These bottom-up approaches simulate individual voter decisions and social interactions, providing a more granular view of electoral dynamics. For instance, Gao et al. (2022a) introduced an agent-based election prediction model that captures social networks and voter-candidate interactions to offer more accurate election forecasts. This approach builds upon earlier work by Lemos et al. (2019), who demonstrated the effectiveness of ABMs in modeling voter turnout and preference formation.
ABMs excel at modeling heterogeneity and dynamic processes in real-world voting scenarios, offering greater flexibility and granularity than traditional statistical models. The work of Collins and Martinez (2023) further illustrates how ABMs can incorporate complex factors such as media influence and peer effects in voter decision-making processes. As the field evolves, future election studies may integrate ABMs with LLMs to balance large-scale data analysis with individual voter behavior modeling, as proposed by Zhang et al. (2024) in their hybrid forecasting framework.
2.3 LLMs in Political Science: Current Research and Our Contributions
@@
In contrast to the existing work, our research provides a more comprehensive approach. In 2024, we introduced an integrated framework that combines LLMs with Agent-Based Models to predict election outcomes and analyze voter behavior. Unlike previous studies, our approach not only focuses on semantic analysis of political texts but also incorporates individual voter behavior modeling, offering more granular and accurate election predictions. This integration addresses the limitations identified by Lopez and Singh (2023) regarding the disconnect between macro-level language models and micro-level voter behavior.
@@ 3.2.2 Version 2: Single-step Prompting with Time-based Information
Capturing macro-level and time-dependent variables is essential for bottom-up agent-based modeling in election prediction Gao et al. (2022b). To enhance the contextual relevance of our simulations, we extended our pipeline by integrating election-year data sourced from Ballotpedia, a well-regarded political information platform. It includes campaign agendas, key policy stances, and candidates’ biographical and professional backgrounds.

### v3

@@ Abstract
Important Notice Ongoing Work: This research is ongoing as of April 4, 2025. Research Integrity: This research is conducted independently w/o any funding. Content Warning: This paper may contain some offensive content generated by LLMs. Disclaimer: This study explores LLMs’ capacity in election forecasting. Predictions do not reflect the authors’ views and should not be interpreted as definitive forecasts.
@@ 2.1 LLMs in Political Science: A New and Emerging Field
The latest work by Li et al. (2024) presents the first systematic framework for integrating LLMs into computational political science. The authors propose a novel taxonomy that categorizes existing work into two main perspectives: political science applications and computational methodologies. From the political science perspective, they highlight LLMs’ capabilities in automating predictive and generative tasks, simulating behavior dynamics, and enhancing causal inference. From the computational perspective, they detail advances in data preparation, fine-tuning approaches, and evaluation methods specifically tailored for political contexts. The paper identifies critical challenges, including the need for domain-specific datasets, addressing bias and fairness issues, incorporating human expertise effectively, and developing evaluation criteria that are aligned with political science requirements. Our work differs from Political-LLM by giving more in-depth analysis on the Decision-Making simulation in Political Science.
@@ 3.2.2 Version 2: Single-step Prompting with Time-based Information
Capturing macro-level and time-dependent variables is essential for bottom-up agent-based modeling in election prediction Gao et al. (2022). To enhance the contextual relevance of our simulations, we extended our pipeline by integrating election-year data sourced from Ballotpedia, a well-regarded political information platform. It includes campaign agendas, key policy stances, and candidates’ biographical and professional backgrounds.